How much do you accumulate if you invest $5,000 @ 8% monthly for 9 years?
Answer by DrBeeee(684)

Please learn the terminology of interest etc. No one will pay you 8% per month! You mean to say the the interest is compounded monthly, much like a mortgage. However the interest rate of 8% is on an annual basis!
The formula you need to use is
(1) B = P(1+r/12)^(12t),
where
B = the balance of your "investment"
P = principal amount invested
r = annual interest rate expressed as a decimal
t = number of years
In your case you should use
(2) B = 5000(1+(.08/12))^(12*9)
Use you calculator to get
(3) B = 5000*2.04995
(4) B = $10,247.65
If youy investment compounded yearly instead of monthly you would get
(5) By = 5000*(1+.08)^9 or
(6) By = $9,995.02 or $252.63 less.